# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v1

@@ 2.3 Analysis of FABind
In this section, we characterize limitations of FABind regarding the important objective in multimodal representation learning, such as intra and shared information. To this end, we consider (3) as
$f^{\rm FB}_{i}={\hbox{arg}}\max_{f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f_{1}({\bf X}_{1});f_{i}({\bf X}_{i})),~{}\forall i\in\{2,\cdots,M\},$ (4)
reflecting the fact that minimizing InfoNCE loss leads to maximizing mutual information. [^3] In contrast to (3), $f_{i}^{\rm FB}$ in (4) might not be aligned with other modalities due to the one-to-one mapping invariant property of mutual information. However, we here do not analyze multimodal alignment of FABind from (4), but rather investigate the performance of encoders in terms of the sufficiency in Definition 1. Let FABind encoders from (4) for each modality be $\mathcal{F}^{\rm FB}=\{f_{1},f_{2}^{\rm FB},\cdots,f_{M}^{\rm FB}\}$. The anchor encoder $f_{1}$ is fixed during the entire FABind procedure. Moreover, we assume that $I(f_{1}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}({\bf X}_{1});{\bf X}_{i})$ is the maximum value that can be achieved by (4) due to data processing inequality (Polyanskiy & Wu, 2024). We next demonstrate that the quality of anchor embedding $f_{1}({\bf X}_{1})$ significantly impacts the performance of $\mathcal{F}^{\rm FB}$ in terms of shared information. The following propositions show the dependency of FABind on anchor embedding quality.
Proposition 1 (FABind with sufficient anchor).
Let $f_{1}^{\rm suf}({\bf X}_{1})$ be a sufficient embedding of the anchor ${\bf X}_{1}$, and let ${\bf X}_{i},i\in[M]$ be discrete. Assume that $f_{i}^{\rm FB},i\in\{2,\cdots,M\}$ are obtained by (4) with a sufficient anchor encoder $f_{1}=f_{1}^{\rm suf}$, i.e., $I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{i})$. Then,
$I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I({\bf X}_{1};{\bf X}_{i}),~{}\forall i\in\{2,\cdots,M\}.$ (5)
@@
The proof can be found in Appendix B.1. ∎
Proposition 2 (FABind with insufficient anchor).
Let $f_{1}^{\rm ins}({\bf X}_{1})$ be insufficient embedding of the anchor ${\bf X}_{1}$ for ${\bf X}_{1}$, in the sense that there exists some $\epsilon>0$ such that $I(f_{1}^{\rm ins}({\bf X}_{1});{\bf X}_{1})<\epsilon\leq\max_{f}I(f({\bf X}_{1});{\bf X}_{1})$. Assume that $f_{i}^{\rm FB},i\in\{2,\cdots,M\}$ are obtained by (4) with $f_{1}=f_{1}^{\rm ins}$, i.e., $I(f_{1}^{\rm ins}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}^{\rm ins}({\bf X}_{1});{\bf X}_{i})$. Then,
@@
Proposition 1 shows that the FABind encoders $\mathcal{F}^{\rm FB}$ learned with a sufficient anchor embedding can achieve the maximum shared information between the anchor and the other modalities. However, it does not guarantee the shared information between non-anchored modalities $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l})),~{}i,l\neq 1$, which can also be seen from (4). Proposition 2 establishes that an insufficient anchor may lead to a reduction of shared information between the anchor and the other modalities, implying that the performance of FABind solely depends on the quality of the anchor.
The analysis reveals several limitations in FABind. Firstly, achieving maximum shared information requires sufficient anchor representation, which depends on having both an informative modality and a sufficient encoder. Without these conditions, FABind may not effectively capture shared information. Secondly, even with sufficient anchor representation, FABind may not provide encoders with maximum intra information. This is because its objective function (4) does not take into account intra information. Thirdly, FABind’s objective function (4) focuses solely on learning shared information between pairs of anchor and non-anchored modalities, while disregarding shared information among non-anchor modalities. This limitation renders FABind less effective when significant shared information exists among non-anchored modalities. Lastly, the representation produced by FABind may not approximate an ideal representation, such as “Platonic representation” (Huh et al., 2024). The Platonic representation is an idealized depiction of reality, which generates all modalities through projections. While integrating all modalities is crucial for constructing such a comprehensive representation, FABind’s approach, which focuses solely on the anchor modality, falls short of this ideal. A more promising direction, which we pursue in this paper, involves learning representations from multiple modalities, fully leveraging fine-grained, sample-level information.
In summary, FABind exhibits the following weaknesses:
  1. P1:
    over-reliance on a single anchor modality;
  2. P2:
    failure to capture intra information;
  3. P3:
    absence of shared information among non-anchored modalities;
  4. P4:
    fundamental limitation of a fixed anchor in representativeness of all modalities.
Next, we propose our method CentroBind, which does not require any choice of anchor modality and is able to capture intra and shared information.
3 Toward a desirable unified representation space
Our method, CentroBind, is a variant of Bind methods, but does not require choosing a single anchor modality. The main intuition deriving CentroBind is as follows: 1) A desirable multimodal embedding should capture representation of all modalities; 2) A desirable unified embedding should attain the highest average alignment. To this end, we generate the anchor representation from all modalities by calculating the centroid of the modalities’ representations. Then, we train the encoders toward minimizing the InfoNCE loss between anchor and other modalities, which is identical to FABind with the generated anchor instead of a fixed anchor. Here we formally describe CentroBind, and show that the method aligns multimodal representations and simultaneously maximizes intra and shared information.
@@ 3 Toward a desirable unified representation space
1:Initialize encoders $f_{1}^{(0)},f_{2}^{(0)},\cdots,f_{M}^{(0)}$.
2:for $t=0,1,\ldots,t_{\max}$ do
3:     Sample a batch dataset $B$ from multimodal datasets $\{\mathcal{D}_{i}\}_{i}$.
4:     Generate anchor embeddings $\{{\boldsymbol{a}}_{j}\}_{j\in\mathcal{I}_{B}}$ using ${\boldsymbol{a}}_{j}={\rm mean}(\{f_{i}^{(t)}({\boldsymbol{x}}_{i,j}^{\prime})\}_{i})$ in (7)
5:     for $i=1,\ldots,M$ do
6:         Optimize $f_{i}^{(t+1)}$ toward minimizing $\mathcal{L}_{\rm CB}(f_{i}^{(t+1)}|\tau)$ in (8)
7:     end for
8:end for
@@ 3.1 CentroBind
To present a general framework for CentroBind, we consider $M$ modalities with corresponding encoders $\{f_{i}\}_{i=1}^{M}$. The algorithmic presentation of CentroBind is in Algorithm 1, and a graphical illustration is given in Figure 1(b). In the following, we elaborate detailed steps of CentroBind.
@@
We initialize $M$ encoders $f_{i}:\mathcal{X}_{i}\to\mathcal{Z},~{}\forall i\in[M]$ for the $M$ modalities. These encoders can either be pretrained models (i.e., backbones) or parameterized models with random weights. The primary constraint for these initial encoders is that their output space must be $\mathcal{Z}$. This ensures that all data samples are mapped to the same output space $\mathcal{Z}$. When using pretrained encoders that produce embeddings in different output spaces, we attach projection layers to standardize the shapes of the embeddings, ensuring consistency of output space across modalities.
@@
Recall that ${\boldsymbol{x}}_{i,j}\in\mathcal{D}_{i}$ denotes the $j$-th feature in the $i$-th modality, where $j$ indexes positive pairs of features (e.g., different views of the same object). In each training iteration of CentroBind, we need to compute an anchor embedding ${\boldsymbol{a}}_{j}$ for the $j$-th multimodal positive features $\{{\boldsymbol{x}}_{i,j}\}_{i=1}^{M}$. This anchor ${\boldsymbol{a}}_{j}$ serves as a desirable aligned embedding for these features. The anchor ${\boldsymbol{a}}_{j}$ is calculated as follows:
@@
where $\text{mean}(\cdot)$ denotes the mean operator that computes the average of its input, and ${\boldsymbol{x}}_{i,j}^{\prime}$ represents an augmented version of ${\boldsymbol{x}}_{i,j}$. If $\{{\boldsymbol{x}}_{i,j}\}_{i=1}^{M}$ are available in multimodal datasets, the anchor is given by ${\boldsymbol{a}}_{j}=\frac{1}{M}\sum_{i=1}^{M}f_{i}({\boldsymbol{x}}_{i,j}^{\prime})$. If only $m<M$ positive pairs are present among $M$ modalities, the anchor is given by ${\boldsymbol{a}}_{j}=\frac{1}{m}\sum_{i\in\mathcal{I}_{j}}f_{i}({\boldsymbol{x}}_{i,j}^{\prime})$, where $\mathcal{I}_{j}$ is the set of indices of modalities having the $m$ available features.
@@
Once anchor embeddings $\{{\boldsymbol{a}}_{j}\}_{j}$ are derived from a batch of data $B=\{{\boldsymbol{x}}_{i,j}\}_{i,j}$, CentroBind aligns each modality-specific encoder’s embedding with the anchor embedding by minimizing the InfoNCE loss. Specifically, let ${\bf A}=\text{mean}(\{f_{i}({\bf X}_{i})\}_{i})$ represent the anchor embedding variable. Then, CentroBind aims to minimize the InfoNCE loss $I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i}))$ across all modalities $i\in[M]$. A detailed expression for this loss is provided in (8).
For symmetry, CentroBind optimizes the following loss function:
@@
$I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i})|\tau)$ $=-\frac{1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})/\tau)}{\sum_{j\in\mathcal{I}_{B}}\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})/\tau)}~{}\text{ and }$ (9a)
$I_{\rm NCE}(f_{i}({\bf X}_{i});{\bf A}|\tau)$ $=-\frac{1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp({\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})/\tau)}{\sum_{j\in\mathcal{I}_{B}}\exp(f_{i}^{\top}({\boldsymbol{x}}_{i,k}){\boldsymbol{a}}_{j}/\tau)}.$ (9b)
@@ 3.2 Theoretical analysis of CentroBind
We start by providing a lower bound of CentroBind’s objective function $\mathcal{L}_{\rm CB}(f_{i}|\tau)$ (8) in Theorem 1, followed by an analysis of minimizing $\mathcal{L}_{\rm CB}(f_{i}|\tau)$.
Theorem 1.
@@
$I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$ $\geq\frac{1}{|\mathcal{I}_{B}|}\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)-\frac{1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i},$ (10)
where $C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ with $g(l,j|k,i):=\exp\left(\frac{|\mathcal{I}_{B}|f_{l}^{\top}({\boldsymbol{x}}_{l,k}^{\prime})f_{i}({\boldsymbol{x}}_{i,j})}{\tau M}\right)$,
$c_{\mathcal{F},k,i}^{\min}=\min_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i),~{}\text{ and }~{}c_{\mathcal{F},k,i}^{\max}=\max_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i).$ (11)
@@
The proof is provided in Appendix B.3. ∎
Theorem 1 provides a lower bound of $I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$ in (9a), which is a part of CentroBind objective $\mathcal{L}_{\rm CB}(f_{i}|\tau)$. Thus CentroBind consequentially minimizes the lower bound (10) that consists of two terms, $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)$ and $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$. We next provide some intuitive explanation on the effect of such a minimization of the lower bound.
The effect of minimizing $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)$.
The objective of minimizing $\sum_{l=1}^{M}I_{\rm NCE}(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\mid\frac{\tau M}{|\mathcal{I}_{B}|})$ involves reducing several InfoNCE losses. Here, each term in the sum represents the InfoNCE loss between embeddings $f_{l}({\bf X}_{l}^{\prime})$ from modality $l$ and $f_{i}({\bf X}_{i})$ from modality $i$, with $\frac{\tau M}{|\mathcal{I}_{B}|}$ being the temperature parameter for scaling the loss. This summation can be divided into two components: 1) Intra Information: When $l=i$, the term measures the similarity between embeddings within the same modality. Minimizing this loss enhances the representation of modality $i$, improving intra information; 2) Shared Information: When $l\neq i$, the term measures the similarity between embeddings from different modalities. Minimizing these losses helps in learning shared information between modalities, contributing to a more coherent multimodal representation.
By optimizing this summation, CentroBind effectively captures both intra and shared information. This results in a more accurate representation for each modality. In contrast, as noted in Section 2.3, FABind does not adequately capture intra information and shared information between non-anchored modalities. This limitation highlights the advantage of CentroBind in achieving a more integrated multimodal representation than fixed anchor binding methods.
@@
We show the effect of growing $C_{\mathcal{F},k,i}$ in terms of cosine similarity score between embeddings. Since $C_{\mathcal{F},k,i}=\frac{1}{4}\left(\sqrt{\gamma}+\sqrt{\frac{1}{\gamma}}\right)^{2}$ with $\gamma=\frac{c_{\mathcal{F},k,i}^{\max}}{c_{\mathcal{F},k,i}^{\min}}\geq 1$, maximizing $C_{\mathcal{F},k,i}$ is equivalent to simultaneously maximizing $c_{\mathcal{F},k,i}^{\max}$ and minimizing $c_{\mathcal{F},k,i}^{\min}$. For ease of the analysis, we assume that the encoders are reasonably well-trained. Then, since a positive pair of embeddings normally yields higher similarity score, $c_{\mathcal{F},k,i}^{\max}$ should be obtained by choosing $l=i$ and $j=k$ in (11) as such choices make ${\boldsymbol{x}}_{l,k}^{\prime}$ be positive pair with ${\boldsymbol{x}}_{i,j}$. Thus, $c_{\mathcal{F},k,i}^{\max}$ is roughly proportional to the similarity score of a positive pair of embeddings. Conversely, $c_{\mathcal{F},k,i}^{\min}$ corresponds to the similarity scores of negative pairs, which tend to be low. Hence, minimizing $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$ enhances the similarity scores for positive pairs and reduces those for negative pairs, improving the overall multimodal alignment.
The preceding analyses demonstrate that CentroBind addresses the limitations $\bf{P1,P2,P3}$, and $\bf{P4}$ of FABind identified in Section 2.3. We now conclude the analysis with a discussion that CentroBind’s unified representation is likely closer to an ideal platonic representation (Huh et al., 2024) compared to FABind’s. A platonic representation is defined as an ideal representation of reality that induces information in all modalities. From this perspective, a representation derived solely from a single modality, without leveraging others, may not be sufficient. In contrast to FABind’s approach, which relies exclusively on the fixed anchor modality, CentroBind constructs its unified representation space by incorporating information from all modalities. This suggests that CentroBind’s unified space is likely to retain a more comprehensive representation of all modalities. Thus, CentroBind emerges as a promising approach for developing a Planotic representation.

### v2

@@ 2.3 Analysis of FABind
In this section, we characterize the theoretical limitations of FABind. To this end, we rewrite (3) as
$f^{\rm FB}_{i}\in{\hbox{arg}}\max_{f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f_{1}({\bf X}_{1});f_{i}({\bf X}_{i})),~{}\forall i\in\{2,\cdots,M\},$ (4)
reflecting the fact that minimizing InfoNCE loss is equivalent to maximizing mutual information. [^3] In contrast to (3), $f_{i}^{\rm FB}$ in (4) might not be aligned with other modalities due to the one-to-one mapping invariant property of mutual information. However, we do not analyze the multi-modal alignment of FABind from (4), but rather investigate the performance of encoders in terms of the sufficiency in Definition 2.1. Let FABind encoders from (4) for each modality be defined as $\mathcal{F}^{\rm FB}=\{f_{1},f_{2}^{\rm FB},\cdots,f_{M}^{\rm FB}\}$. The anchor encoder $f_{1}$ is fixed during the entire FABind procedure. Moreover, we assume that $I(f_{1}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}({\bf X}_{1});{\bf X}_{i})$ is the maximum value that can be achieved by (4) due to data processing inequality (Polyanskiy & Wu, 2024). We next demonstrate that the quality of anchor embedding $f_{1}({\bf X}_{1})$ significantly impacts the performance of $\mathcal{F}^{\rm FB}$ in terms of shared information. The following propositions show the dependency of FABind on anchor embedding quality.
Proposition 2.2 (FABind with sufficient anchor).
Let $f_{1}^{\rm suf}({\bf X}_{1})$ be a sufficient embedding of the anchor ${\bf X}_{1}$, and let ${\bf X}_{i},i\in[M],$ be a discrete random variable. Assume that $f_{i}^{\rm FB},i\in\{2,\cdots,M\}$ are obtained by (4) with a sufficient anchor encoder $f_{1}=f_{1}^{\rm suf}$, i.e., $I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{i})$. Then, for all $i\in\{2,\cdots,M\}$,
$I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I({\bf X}_{1};{\bf X}_{i}).$ (5)
@@
The proof is in Appendix B.1. ∎
Proposition 2.3 (FABind with insufficient anchor).
Let $f_{1}^{\rm ins}({\bf X}_{1})$ be an insufficient embedding of the anchor ${\bf X}_{1}$ for ${\bf X}_{1}$, in the sense that there exists some $\epsilon>0$ such that $I(f_{1}^{\rm ins}({\bf X}_{1});{\bf X}_{1})<\epsilon\leq\max_{f}I(f({\bf X}_{1});{\bf X}_{1})$. Assume that $f_{i}^{\rm FB},i\in\{2,\cdots,M\}$ are obtained by (4) with $f_{1}=f_{1}^{\rm ins}$, i.e., $I(f_{1}^{\rm ins}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i}))=I(f_{1}^{\rm ins}({\bf X}_{1});{\bf X}_{i})$. Then,
@@
Proposition 2.2 shows that the FABind encoders $\mathcal{F}^{\rm FB}$ learned with a sufficient anchor embedding can achieve the maximum shared information between the anchor and the other modalities. However, it does not guarantee shared information between non-anchored modalities $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l})),~{}i,l\neq 1$, which can also be seen from (4). Proposition 2.3 establishes that an insufficient anchor may lead to a reduction of shared information between the anchor and the other modalities, implying that the performance of FABind may overly depend on the quality of the arbitrarily selected anchor. Next, we enumerate three limitations in FABind:
P1. Over-reliance on a single anchor modality.
Achieving maximum shared information requires sufficient anchor representation (Proposition 2.2 and 2.3), which depends on having both an informative modality and a sufficient encoder. Without these conditions, FABind may not effectively capture shared information.
P2. Failure to capture intra information.
Even with sufficient anchor representation, FABind may not provide encoders with maximum intra information. This is because its objective function (4) does not take into account the mutual information between a sample from a modality and its augmentation, i.e., $I(f_{i}({\bf X}_{i});{\bf X}_{i})$.
P3. Absence of shared information among non-anchored modalities.
The objective function of FABind (4) focuses solely on learning shared information between anchor and non-anchored modalities, while disregarding shared information among non-anchored modalities. This implies that FABind may not capture shared information among them. This limitation could render FABind less effective when crucial shared information exists among non-anchored modalities.
We identify three key limitations (P1, P2, and P3) of FABind, supported by empirical evidence presented in Section 4.1. In addition to these constraints, the representations generated by FABind may not approximate an ideal standard, such as the “Platonic representation” described by Huh et al. (2024). Achieving an optimal multi-modal representation necessitates the comprehensive integration of all modalities to fully capture their information. However, FABind falls short of this objective. To address these shortcomings, we next introduce an alternative multi-modal representation that fully leverages sample-level information across all modalities.
3 Alignment using adaptive anchor
To address the limitations of fixed modality anchoring, we propose the concept of adaptive anchor alignment. We aim for a unified representation that (1) remains unbiased toward any single modality and (2) achieves high similarity alignment across all modalities. In this paper, we introduce a centroid-based anchor representation that naturally fulfills these two requirements by acting as a geometric center. Specifically, we train the encoders by minimizing the InfoNCE loss between each modality and the centroid (similar to FABind). Although other adaptive anchor methods—such as median or weighted average—are viable alternatives (see Appendix C.1), we focus on the centroid for its simplicity and strong theoretical grounding in aligning multi-modal representations. Next, we formally define CentroBind and show how it maximizes both intra-modal and cross-modal information within a unified embedding space.
@@ 3 Alignment using adaptive anchor
  Initialize encoders $f_{1}^{(0)},f_{2}^{(0)},\cdots,f_{M}^{(0)}$.
  for $t=0,1,\ldots,t_{\max}$ do
     Sample $B$ from multi-modal datasets $\{\mathcal{D}_{i}\}_{i}$.
     Generate anchor embeddings $\{{\boldsymbol{a}}_{j}\}_{j\in\mathcal{I}_{B}}$ using (7)
     for $i=1,\ldots,M$ do
        Minimize $\mathcal{L}_{\rm CB}(f_{i}^{(t+1)}|\tau)$ in (8)
     end for
  end for
@@ 3.1 CentroBind
Consider $M$ modalities with corresponding encoders $\{f_{i}\}_{i=1}^{M}$. The CentroBind algorithm is presented in Algorithm 1, and a graphical illustration is given in Figure 2. In the following, we describe each step of the algorithm.
@@
We initialize $M$ encoders $f_{i}:\mathcal{X}_{i}\to\mathcal{Z},~{}\forall i\in[M]$ for the $M$ modalities. These encoders can either be pretrained models (i.e., backbones) or parameterized models with random weights. The primary constraint for these initial encoders is that their output space must be the modality-independent embedding space $\mathcal{Z}$. When using pretrained encoders that produce embeddings in different output spaces, these are projected onto the embedding space $\mathcal{Z}$, ensuring consistency of output space across modalities.
@@
Recall that ${\boldsymbol{x}}_{i,j}\in\mathcal{D}_{i}$ denotes the $j$-th feature in the $i$-th modality, where $j$ indexes positive pairs of features (e.g., different views of the same object). In each training iteration of CentroBind, we need to compute an anchor embedding ${\boldsymbol{a}}_{j}$ for the $j$-th multi-modal positive features $\{{\boldsymbol{x}}_{i,j}\}_{i=1}^{M}$. This anchor ${\boldsymbol{a}}_{j}$ serves as a desirable aligned embedding for these features. The anchor ${\boldsymbol{a}}_{j}$ is calculated as follows:
@@
where $\text{mean}(\cdot)$ denotes the mean operator that computes the average of its input, and ${\boldsymbol{x}}_{i,j}^{\prime}$ represents an augmented version of ${\boldsymbol{x}}_{i,j}$. If $\{{\boldsymbol{x}}_{i,j}\}_{i=1}^{M}$ are available in multi-modal datasets, the anchor is given by ${\boldsymbol{a}}_{j}=\frac{1}{M}\sum_{i=1}^{M}f_{i}({\boldsymbol{x}}_{i,j}^{\prime})$. If only $m<M$ positive pairs are present among $M$ modalities, the anchor is given by ${\boldsymbol{a}}_{j}=\frac{1}{m}\sum_{i\in\mathcal{I}_{j}}f_{i}({\boldsymbol{x}}_{i,j}^{\prime})$, where $\mathcal{I}_{j}$ is the set of indices of modalities having the $m$ available features.
@@
Once anchor embeddings $\{{\boldsymbol{a}}_{j}\}_{j}$ are derived from a batch $B=\{{\boldsymbol{x}}_{i,j}\}_{i,j}$, CentroBind aligns each modality-specific encoder embedding with the anchor embedding by minimizing InfoNCE loss. Specifically, let ${\bf A}=\text{mean}(\{f_{i}({\bf X}_{i})\}_{i})$ represent the anchor embedding variable. Then, CentroBind aims to minimize InfoNCE loss $I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i}))$ across all modalities $i\in[M]$. A detailed expression for this loss is provided in (8).
CentroBind optimizes the following symmetrized loss function:
@@
$I_{\rm NCE}({\bf A};f_{i}({\bf X}_{i})|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})}{\tau})}{\sum_{j\in\mathcal{I}_{B}}\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})}{\tau})}$ (9a)
$I_{\rm NCE}(f_{i}({\bf X}_{i});{\bf A}|\tau)$ $=\frac{-1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log\frac{\exp(\frac{{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})}{\tau})}{\sum_{j\in\mathcal{I}_{B}}\exp(\frac{f_{i}^{\top}({\boldsymbol{x}}_{i,k}){\boldsymbol{a}}_{j}}{\tau})}.$ (9b)
@@ 3.2 Theoretical analysis of CentroBind
We start by providing a lower bound on the objective function of CentroBind  $\mathcal{L}_{\rm CB}(f_{i}|\tau)$ (8) in Theorem 3.1, followed by an analysis of the minimizer of $\mathcal{L}_{\rm CB}(f_{i}|\tau)$.
Theorem 3.1.
@@
$|\mathcal{I}_{B}|I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$
$\geq\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i},$ (10)
where
$C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ (11)
with $g(l,j|k,i):=\exp\left(\frac{|\mathcal{I}_{B}|f_{l}^{\top}({\boldsymbol{x}}_{l,k}^{\prime})f_{i}({\boldsymbol{x}}_{i,j})}{\tau M}\right)$,
$c_{\mathcal{F},k,i}^{\min}=\min_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i),~{}\text{ and }~{}$
$c_{\mathcal{F},k,i}^{\max}=\max_{l\in[M],j\in\mathcal{I}_{B}}g(l,j|k,i).$ (12)
@@
The proof is in Appendix B.3. ∎
Theorem 3.1 provides a lower bound of $I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;{\tau}\right)$ in (9a), which is a part of the CentroBind objective function $\mathcal{L}_{\rm CB}(f_{i}|\tau)$. Thus CentroBind minimizes a lower bound (3.1) that consists of two terms, $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{{\tau}M}{|\mathcal{I}_{B}|}\right)$ and $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$. We next provide intuition on why a minimization of the lower bound is justified.
The effect of minimizing $\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\cdot\right)$.
The objective of minimizing $\sum_{l=1}^{M}I_{\rm NCE}(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\mid\frac{\tau M}{|\mathcal{I}_{B}|})$ is to reduce several InfoNCE losses. Here, each term in the sum represents InfoNCE loss between embeddings $f_{l}({\bf X}_{l}^{\prime})$ from modality $l$ and $f_{i}({\bf X}_{i})$ from modality $i$, with $\frac{\tau M}{|\mathcal{I}_{B}|}$ being a temperature parameter for scaling the loss. This summation can be divided into two components: 1) Intra Information: When $l=i$, the term measures the similarity between embeddings within the same modality. Minimizing this loss enhances the representation of modality $i$, improving intra information; 2) Shared Information: When $l\neq i$, the term measures the similarity between embeddings from different modalities. Minimizing these losses helps in learning shared information between modalities, contributing to a more representative multi-modal embedding.
By optimizing this summation, CentroBind effectively captures both intra and shared information. As shown below, this generally results in a more balanced representation for the modalities. In contrast, as noted in Section 2.3, FABind does not adequately capture intra information and shared information between non-anchored modalities. This limitation highlights the advantage of CentroBind in achieving a more integrated multi-modal representation than fixed anchor binding methods.
@@
We show the effect of growing $C_{\mathcal{F},k,i}$ in terms of cosine similarity score between embeddings. Since $C_{\mathcal{F},k,i}=\frac{1}{4}\left(\sqrt{\gamma}+\sqrt{\frac{1}{\gamma}}\right)^{2}$ with $\gamma=\frac{c_{\mathcal{F},k,i}^{\max}}{c_{\mathcal{F},k,i}^{\min}}\geq 1$, maximizing $C_{\mathcal{F},k,i}$ is equivalent to simultaneously maximizing $c_{\mathcal{F},k,i}^{\max}$ and minimizing $c_{\mathcal{F},k,i}^{\min}$. For ease of analysis, we assume that the encoders are reasonably well-trained. Then, since a positive pair of embeddings normally yields higher similarity score, $c_{\mathcal{F},k,i}^{\max}$ is attained by choosing $l=i$ and $j=k$ in (3.1) as such choices make ${\boldsymbol{x}}_{l,k}^{\prime}$ be positive pair with ${\boldsymbol{x}}_{i,j}$. Thus, $c_{\mathcal{F},k,i}^{\max}$ is roughly proportional to the similarity score of a positive pair of embeddings. Conversely, $c_{\mathcal{F},k,i}^{\min}$ corresponds to the similarity scores of negative pairs, which tend to be low. Hence, minimizing $-\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i}$ enhances the similarity scores for positive pairs and reduces those for negative pairs, improving the overall multi-modal alignment.
These comments suggest that CentroBind addresses the limitations $\bf{P1,P2}$, and $\bf{P3}$ of FABind identified in Section 2.3.